OpenStudy (anonymous):

A jogger runs 300 m in the fi rst minute. The distance the jogger covers decreases by 20 m in each succeeding minute. What distance does the jogger cover in the 7th minute? How do I come up with an equation?

OpenStudy (anonymous):

making a table can help: min distance 1 300m 2 300-20 3 300-20-20 4 300-20-20-20 we can see that at 2 min we subtract one twenty at 3 min we subtract two twenties at 4 min we subtract three twenties. total distance = 300m - 20*minutes beyond 1 minute becomes total distance = 300m - 20(min - 1) or distance = 300 - 20(t-1); where t is time.

OpenStudy (anonymous):

but doesn't that mean that the jogger covers 280m in the second minute? shouldn't that be 580m?

OpenStudy (anonymous):

shoot yes, you are right, sorry I did the arithmetic term, to get the sum, we have to use the arithetric sum formula, so, d =(t/2)*(2*300-20*(t-1)) sorry about that

OpenStudy (anonymous):

thanks for your help!